Q6. There 40 students in a chemistry class and 60 students in physics class. Find the number of students which are either in Physics class or Chemistry class in the cases. (i) the two classes meet at the same hour. (ii) the two classes meet at different hours and 20 students are enrolled in both the subjects.  
can u pls explain clearly im not able to understand the question
 

Grade:12

1 Answers

Vikas TU
7 years ago
Total students when they meet at the same hour.
60 + 40 = 100.
In this 100 students the common would be:
60 – 40 = 20.
the AUB = A + B – AintersectionB = > 60 + 40 – 20 = > 80